OpenStudy (anonymous):

Find the equation of the line that passes through the points (1, 2) and (-2, 5)

OpenStudy (anonymous):

find the slope

OpenStudy (anonymous):

y = (-7/3)x + 1?

OpenStudy (anonymous):

the slope is (5-2)/(-2-1) = -1 then the equation becomes y=-x+c, c is 3 y=-x+1

OpenStudy (anonymous):

I think I was right

OpenStudy (anonymous):

SLOPE = -1 Y INTERCEPT = 3

OpenStudy (anonymous):

@Lib Libs you did wrong calculation for the slope

OpenStudy (anonymous):

y = 3x + 3 y = (-7/3)x + 1 y = x + 3 y = -x + 3 It seemed like the best out of these choices

OpenStudy (anonymous):

THAT'S IT

OpenStudy (anonymous):

So 3x +3?

OpenStudy (anonymous):

Riight?

OpenStudy (anonymous):

y = -x + 3